A helium-neon laser emits light of wavelength 632.8 nm and has a power output of 2.0 mW. How many photons are emitted per second by this laser?
________ photons/s
Answer #1

E in photon = hc/lambda = (6.62E-34) * (3E8) / 632.8E-9 = 3.13843E-19 J

Photons per second = Total Power / E_photon = 3E-3 / 3.13843E-19 = 9.55892E+15 photons/second
Answer #2

Energy of a photon at 632.8 nm is

E_photon = hc/lambda = (6.62E-34) * (3E8) / 632.8E-9 = 3.13843E-19 J

Photons per second = Total Power / E_photon = 2E-3 / 3.13843E-19 = 6.37261E+15 photons/second